kg) have the same momentum. At what speed is the ball 9. A photon of red light (wavelength = 730 nm) and a Ping-Pong ball (mahi please show your work clearly and explain, thank you!
Answer #1

momentum of photon, p = (h / lambda) = (6.626 X10-34) /(730 X10-9) = 9.0768X10-28 kg.m/s

speed of the ball = (p / m) = (9.0768X10-28 ) / (2.30 X10-3) = 3.9464 X10-25 m/s (Ans)